# Universal Omni-Language Models

### v4

@@ 5.2 The Effectiveness of OmniInstruct
5.3 Textual Approximation on Images and Audios
As the absence of strong OLM baselines, we further introduce the text alternatives of images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal MLLMs to analyze the current research progress in the field.
Performance Changes of Open OLMs.
Performances of Dual-modal MLLMs.
In the setting of using text as the alternatives of audios and images, the VLMs show generally better results than ALMs (Table 6 vs Table 7) even compared with open-source model with similar model size [^4] The results of pure text $(I^{\prime},A^{\prime})$ setting are placed at Table 8, Appendix A.. This could be caused by : 1) more available research resources have been put in VLMs to develop datasets and cross-modality alignment architectures, leading to higher instruction following rate and accuracy compared to ALMs; 2) the audio data are naturally harder (and hence more expensive) to annotate; and 3) audio typically has longer sequence tokens and requires more computational resource compared to text and image, making it harder to scale up. If $I^{\prime}$ and $A^{\prime}$ have the information loss ratio when converted from $I$ and $A$, it seems to be easier for the researchers to train the future omni-language models from exisiting VLMs rather than ALMs. Besides, we can observe Claude-3.5 and GPT-4o are generally the best two VLMs, significantly better compared to open-source VLMs. Qwen2- audio and Gemini are the two best ALMs in speech, Qwen2- audio is the best in sound, and Audio-SALMONN is music. Moreover, we can see significant differences in different types of audio, i.e., LTU and audio-flamingo are worse for music compared to speech, while Qwen-audio, which includes music on pre-training, provides better results on music compared to speech. And MusiLingo only uses music for pre-training performs worse in speech and audio.
@@ 5.3 Textual Approximation on Images and Audios
The performance gaps brought by the replaced textual image and audio descriptions are in revealed in Figure 4. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models show stronger reasoning capability when equipped with information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. But the gains of best proprietary models GPT4-o (0513) andClaude-3.5-Sonnet exhibit is not significant, though GPT4-o (0513) achieved an impressive 60.60% accuracy in the pure textual setting.
@@ Appendix A More Experiment Results
Table 8: Results on Pure Textual Approximation for Both Image and Audio. All the images and audios are represented in texts. The results at the second and third column are taken from the corresponding models in Table 6 and Table 7.
@@ B.1 Statistics for OmniInstruct Dataset
B.2 Prompt for Quality Control on OmniInstruct Dataset
Figure 5: The Prompt for OmniInstruct Dataset Filtering.
@@ B.3 Diversity of Music Audios of OmniBenchmark
Appendix C Statistic Significance of the Findings
@@
C.1 Inference if Two Models Provide Significantly Different Results under the Same Small Subset
The method is inspired by  (Guyon et al., 1998). Suppose all the data samples $\{x_{i}\}_{i=1}^{n}$ are i.id, where $n$ is the number of data samples. For each specific model $M_{j}$, the performance of the $j^{th}$ model on the $i^{th}$ data is $M_{j}\left(x_{i}\right)$ which is equal to 00 if the model predicts the wrong choice and $1$ otherwise. Due to the i.id. assumption of the test set, the $\{M_{j}\left(x_{i}\right)\}_{i=1}^{n}$ is a random variable sequence (r.v.s.) with binomial distribution with mean $p_{i}$ and variance $\sigma_{i}^{2}=p_{i}(1-p_{i})$.
@@ C.1 Inference if Two Models Provide Significantly Different Results under the Same Small Subset
SImilarly, with Chernoff bound  (Chernoff, 1952), $p-\hat{p}\leq\sqrt{\frac{-2p\left(\ln\alpha\right)}{n}}$ hold with the probability $1-\alpha$. And therefore, $n=\frac{-2\ln\alpha}{\beta^{2}p}$. The alternative estimation of p-value is $\alpha=\exp\left(\frac{np\beta^{2}}{2}\right)$
@@
    • Claim (6): **In speech subset (771 samples)** demonstrated in table 3 with audio and image as input, Gemini-1.5-Pro (42.67%) does not surpass UnifiedIO2-xlarge (3.2B) (39.56%) significantly, but it surpass all other models like the second-best UnifiedIO2-xxlarge (6.8B) (34.24%) with p-value less than 0.01
    • Claim (7): **In general audio subset (265 samples)** demonstrated in Table 3 with audio and image as input, Gemini-1.5-Pro (42.26%) is better than all the others besides UnifiedIO2, such as Video-SALMONN (13B) (31.70%) with p-value less than 0.01
    • Claim (8): **In the music subset (106 samples)** demonstrated in Table 3 with audio and image as input, Video-SALMONN (13B) (56.6%) is better than all the other models such as the best one Gemini-1.5-Pro (46.23%) with p-value less than 0.05.
@@
C.2 Comparasion between Two Different Setting with the Same Samples
@@
Recaall the performance of UnifiedIO2 in different experimental settings are as follows:
@@
  • Claim (17): There is no significant difference between Table 2 and Table 5 (p-value 0.562) or Table 6 (p-value 0.919), showing the capability of OLM on image understanding is similar to the SOTA LLM generated image caption.
C.3 Comparasion on the Same Setting with Two Different Subset Samples
@@
Given that we conducted 20-30 hypothesis tests, it is not surprising to observe at least one p-value below 0.05 due to the multiple comparisons, which could suggest a false positive in some cases. However, many of the tests involve the same models or subsets, meaning the p-values are not independent. As a result, the observed number of significant p-values (¡ 0.04, in fact, ¡ 0.01 for many cases) is less indicative of false positives than it would be under the assumption of independent tests. This reduces the likelihood of false positives in our findings.
Appendix D Human Evaluation

### v6

@@ 5.3 Textual Approximation on Images and Audios
The performance gaps brought by the replaced textual image and audio descriptions are in revealed in Figure 6. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models show stronger reasoning capability when equipped with information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. But the gains of best proprietary models GPT4-o (0513) andClaude-3.5-Sonnet exhibit is not significant, though former achieved an impressive 60.60% in such setting.
@@ Appendix A More Experiment Results
Table 6: Results on Pure Textual Approximation for Both Image and Audio. All the images and audios are represented in texts. The results at the second and third column are taken from the corresponding models in Figure 5 and Table 5.
@@ B.1 Statistics for OmniInstruct Dataset
B.2 Discussion of the Human Annotation Protocol
Question creation involved 16 annotators and 5 quality reviewers, covering music, linguistics, phonetics, cryptography, and math. During verification, annotators were required to document the modality-specific rationale for each question. Verifiers explicitly assessed this rationale for consistency. This supports our claim that questions are difficult, but not ill-defined. The final human test please refer to Appendix E.
@@ B.3 Diversity of Music Audios of OmniBenchmark
B.4 The Discussion of Video Data
While video data naturally encodes richer temporal dynamics, our current focus was to study targeted cross-modal reasoning using static images and audio. This choice enables finer control over modality composition and ambiguity—images can be selectively edited, cropped, or paired with audio to emphasize specific reasoning tasks (e.g., conflicting cues or missing information). At the time of benchmark development, only GPT-4o and Gemini offered reliable video+audio capabilities, whereas most open-source video models lacked support for audio streams, making it difficult for the current version of benchmark to include video-audio reasoning. Moreover, we believe that pruning the complexity of the temporal vision information could be a clearer experimental setting to analyze whether an OLM can better understand and reason based on the cross vision-audio information.
B.5 Prompt for Quality Control on OmniInstruct Dataset
Figure 7: The Prompt for OmniInstruct Dataset Filtering.
B.6 Audio Transcription and Caption
We propose image caption as well as audio transcription and caption results to test ALMs and VLMs. For speech and song, we transcribe the speech text or lyrics. For music and other audio events that lacks natural language transcripts (e.g., instrumental music), we provide expert-written annotations capturing salient audio features such as instrumentation, emotion, dynamics, texture, genre, and production characteristics. These annotations are paired with the image and text to form the necessary context to answer the question. All music-based questions in OmniBench were manually vetted to ensure answerability using image and audio transcriptions alone.
We did not experiment with spectrogram-based embeddings for VLMs, though this maybe a promising direction. However, several technical limitations motivated our choice to use expert-written audio descriptions. On the one hand, spectrogram type and resolution vary by task: Speaker ID often requires 75–120 ms STFT windows; ASR uses 20–35 ms; pitch detection prefers CQT. A single universal spectrogram representation does not exist. On the other hand, interpretability: Expert-written audio summaries (e.g., instrumentation, emotion, genre, acoustics) paired with image and text offer a semantically compact, modality-aligned representation accessible to current vision-language models. Nevertheless, we highlight this as a direction for extending VLM compatibility with audio tasks.
Appendix C Data Annotation Platform
We conducted annotation correction and quality inspection using a professional annotation platform that supports structured editing, version control, and multi-stage review. A Snapshot of the platform interface is shown in Figure 8.
Figure 8: A Snapshot of the data annotation platform for human evaluation.
Appendix D Statistic Significance of the Findings
@@
D.1 Inference if Two Models Provide Significantly Different Results under the Same Small Subset
The method is inspired by  [Guyon et al., 1998]. Suppose all the data samples $\{x_{i}\}_{i=1}^{n}$ are i.id, where $n$ is the number of data samples. For each specific model $M_{j}$, the performance of the $j^{th}$ model on the $i^{th}$ data is $M_{j}\left(x_{i}\right)$ which is equal to 0 if the model predicts the wrong choice and $1$ otherwise. Due to the i.id. assumption of the test set, the $\{M_{j}\left(x_{i}\right)\}_{i=1}^{n}$ is a random variable sequence (r.v.s.) with binomial distribution with mean $p_{i}$ and variance $\sigma_{i}^{2}=p_{i}(1-p_{i})$.
@@ D.1 Inference if Two Models Provide Significantly Different Results under the Same Small Subset
SImilarly, with Chernoff bound  [Chernoff, 1952], $p-\hat{p}\leq\sqrt{\frac{-2p\left(\ln\alpha\right)}{n}}$ hold with the probability $1-\alpha$. And therefore, $n=\frac{-2\ln\alpha}{\beta^{2}p}$. The alternative estimation of p-value is $\alpha=\exp\left(\frac{np\beta^{2}}{2}\right)$
@@
    • Claim (6): **In speech subset (771 samples)** demonstrated in Figure 4 with audio and image as input, Gemini-1.5-Pro (42.67%) does not surpass UnifiedIO2-xlarge (3.2B) (39.56%) significantly, but it surpass all other models like the second-best UnifiedIO2-xxlarge (6.8B) (34.24%) with p-value less than 0.01
    • Claim (7): **In general audio subset (265 samples)** demonstrated in Figure 4 with audio and image as input, Gemini-1.5-Pro (42.26%) is better than all the others besides UnifiedIO2, such as Video-SALMONN (13B) (31.70%) with p-value less than 0.01
    • Claim (8): **In the music subset (106 samples)** demonstrated in Figure 4 with audio and image as input, Video-SALMONN (13B) (56.6%) is better than all the other models such as the best one Gemini-1.5-Pro (46.23%) with p-value less than 0.05.
  • Compared with Table 5
@@
D.2 Comparasion between Two Different Setting with the Same Samples
@@
Recall the performance of UnifiedIO2 in different experimental settings are as follows:
@@
  • Claim (17): There is no significant difference between Table 3 and Figure 5 (p-value 0.562) or Table 5 (p-value 0.919), showing the capability of OLM on image understanding is similar to the SOTA LLM generated image caption.
D.3 Comparasion on the Same Setting with Two Different Subset Samples
@@
Given that we conducted 20-30 hypothesis tests, it is not surprising to observe at least one p-value below 0.05 due to the multiple comparisons, which could suggest a false positive in some cases. However, many of the tests involve the same models or subsets, meaning the p-values are not independent. As a result, the observed number of significant p-values (< 0.04, in fact, < 0.01 for many cases) is less indicative of false positives than it would be under the assumption of independent tests. This reduces the likelihood of false positives in our findings.
Appendix E Human Evaluation
All the questions is reviewed by evaluators (denoted as “human evaluator” or “human expert” in the main text) with X-minute answer time limit (presented in the draft) and another set of experts with more time permit (2-10 minutes), deriving a higher score of 74.03%. This includes the comparison between humans with strict and loose time limits to acquire responses to estimate difficulty vs. ambiguity.
@@ Appendix E Human Evaluation
Appendix F Limitation
Despite the strengths of OmniBench, several limitations remain. First, due to the high cost of annotation—each sample requiring approximately 30 minutes to design and 10 minutes to verify by human annotators—this version of the benchmark includes only 1,000 manually curated examples. We plan to expand the dataset in future iterations. Second, the image captions used for text-alternative settings were generated by large language models rather than human annotators. While this approach improves scalability, it leads to suboptimal performance on tasks requiring precise visual text recognition (e.g., OCR-related questions), and the caption quality may not consistently surpass human-generated descriptions. Lastly, the current version of OmniBench focuses on static images and does not yet include video inputs. Extending the benchmark to support temporal reasoning with video and audio remains an important direction for future work.
Appendix G Experiments Compute Resources
The inference experiments were conducted using A800 GPUs, with each model processing the full OmniBench dataset over a period ranging from 30 minutes to 3 hours, depending on the model size and complexity. The evaluations were conducted via public APIs, which introduced variability due to network latency. The supervised fine-tuning on OmniInstruct dataset takes 5 minutes with LoRA using eight H800 GPUs for each epoch.
Appendix H Ethical Statement
Human annotation and fair wages: All annotators involved in data creation were either legally employed research assistants with music professionals or students supported by formal scholarships who are all co-authors of the paper, and were compensated in accordance with local minimum wage regulations. This aligns with NeurIPS requirements for fair compensation of human participants.
Data privacy and consent: All audio clips in MMAR were extracted from publicly accessible, user-uploaded videos on platforms like YouTube. Each clip is shorter than 30 seconds—shorter than preview segments on commercial platforms like Spotify—minimizing copyright and privacy risks. No personally identifiable or sensitive user information is included.
Licensing and responsible use: The dataset will be released under a CC-BY-NC license, explicitly limiting its use to non-commercial academic research, in accordance with NeurIPS guidelines for ethical dataset release and copyright respect.
Diversity and representativeness: OmniBench includes a balanced range of speech and vocal music data as discussed in subsection B.3, and the dataset covers multiple spoken and sung languages, including English, Chinese, Japanese, French, Italian, Spanish and Russian etc. This reflects an effort toward diversity and mitigates representational bias.
Environmental and safety considerations: The research poses no foreseeable risks related to safety, security, discrimination, surveillance, or environmental harm. The benchmark does not involve high-compute training or deployment of risky generative models.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the abstract and introduction accurately reflect the paper’s contributions and scope. The introduction of OmniBench and OmniInstruct aligns with the stated goal of evaluating and improving tri-modal reasoning in MLLMs. The main findings—performance limitations of open-source models, fragility of proprietary models, and human-model performance gaps—are consistent with the claims made. Overall, the abstract provides a clear and faithful summary of the paper’s objectives, methods, and insights.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation is discussed in Appendix F. These reflections demonstrate awareness of the benchmark’s scope and encourage future improvement.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not present any theoretical results requiring formal assumptions or mathematical proofs. The only formulaic reference is the multiple methods on hypothesis testing, which is a standard statistical adjustment rather than a novel theoretical contribution. We provide the details calculation in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides a comprehensive and transparent account of all components necessary to reproduce its main experimental results. The OmniBench benchmark and OmniInstruct dataset, including both JSON annotations and audio files, is publicly hosted on Hugging Face, and the evaluation script is available on GitHub. Detailed descriptions are provided for all evaluated models that are either open-sourced or have APIS available, including citations and categorisation across different input modalities. The paper also clearly explains the evaluation procedure, including how image captions are generated and used as inputs to non-visual models. This level of detail ensures reproducibility of both the experimental setup and the main benchmarking results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper provides open access to both the dataset and evaluation code. The OmniBench benchmark and OmniInstruct dataset, including audio files and JSON annotations, is released on Hugging Face, while the evaluation script and usage instructions are hosted on GitHub.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies all relevant experimental details necessary to understand and reproduce the results. The dataset curation process is extensively documented, including how questions, answers, and image/audio transcription were created, validated, and quality-checked. The evaluation setting clearly defines the task formulation (multiple-choice), metric (classification accuracy), and model input structure (audio + question + choices), along with the curation of OmniInstruct dataset. For ALM and VLM, the paper describes how image/audio is first converted to captions before being fed into language models. Although the paper does not involve model training and therefore omits hyperparameter or optimizer details, it thoroughly covers all required test-time settings. Model categories, individual model sources, and evaluation procedures are well-documented.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper provides a detailed description of three different types of hypothesis testing in Appendix D.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Information discussed in Appendix G. While exact memory and storage configurations are not specified for each run, the provided information is sufficient to estimate the compute scale and replicate the experimental setup.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms to the NeurIPS Code of Ethics in all respects, including ethical data sourcing, fair labor practices, legal compliance, privacy protection, and responsible dissemination in Appendix H By adhering to these ethical standards throughout data collection, annotation, and release, the work aligns with the spirit and letter of the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The paper presents a benchmark dataset for academic research and does not involve the development or deployment of models with direct real-world applications. It does not include personal data, sensitive content, or generative components, and all audio clips are sourced from publicly available material and limited to under 30 seconds. Given its foundational nature and restricted non-commercial license (CC-BY-NC), the work does not pose immediate societal impact risks.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release any models or data with high risk for misuse. The OmniBench benchmark and OmniInstruct dataset contains short (<30s) audio clips sourced from publicly available videos, with no personal or sensitive content. It does not involve the release of generative models or systems that could be repurposed for harmful applications. As such, safeguards for high-risk assets are not necessary in this context.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All existing assets used in OmniBench, including pre-trained models and datasets, are properly cited with references to their original publications. The terms of use for each asset are respected, and our benchmark is released under the MIT license. Audio data is sourced from publicly available, user-uploaded internet content, and is used under fair-use considerations with additional restrictions (e.g., <30 seconds, CC-BY-NC license) to minimize copyright risks. The OmniInstruct dataset is based on AVQA (non-commercial license), MSRVTT-QA (MIT license), and MUSIC-AVQA-v2.0 (license), and therefore, we setup the OmniInstruct as cc-by-nc license as well.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The paper introduces a new benchmark (OmniBench) and a new dataset (OmniInstruct). all associated assets—including audio files, annotations, and evaluation scripts—are well documented in the README files of both the Hugging Face dataset repository and the GitHub codebase.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: The paper does not involve crowdsourcing or external human subject experiments. All annotations were performed by the authors themselves or employees from the same company as the correspondent authors, being paid much more than the minimum wage in the region of the data collector. The process of annotation and human evaluation is discussed in the paper. To ensure transparency, the paper includes screenshots of the annotation interface in Appendix C.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: All data collection and annotation were conducted by the authors and qualified collaborators or the employees of the company. This study has been reviewed and approved by the Human and Artefacts Research Ethics Committee of legal team at 01.ai, ensuring our research adheres to ethical guidelines in data usage, AI generation, and cultural representation. No direct interaction with participants occurred, and no personally identifiable information was collected.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLMS were used to determine whether the multiple-choice options of each question are confusing enough. Besides, the image caption is generated by VLM to evaluate the performance of ALM. Moreover, VLM is used to generate and filter the data quality of the OmniInstruct dataset.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.